A free-fall ride takes 2.35 s. to fall from it’s maximum height. How tall is the ride?
1
Expert's answer
2019-06-06T10:00:39-0400

For free fall we write [1]


"y=y_0+v_0t-\frac{gt^2}{2}"

We have "t=2.35\,s, \, y(2.35)=0,\, v_0=0, g=9.8\, m/s^2." So,

"y_0=\frac{gt^2}{2}=\frac{9.8 \cdot 2.35^2}{2}\approx 27.06\, m."

[1] https://en.wikipedia.org/wiki/Free_fall#Free_fall_in_Newtonian_mechanics